mhchen:

arclength of ln(sinx) from pi/4 to 3pi/4:

mhchen:

So I use the arc-length formula: \[\int\limits_{\pi/4}^{3\pi/4}\sqrt{1+(\frac{d}{dx}\ln(\sin(x)))^2}dx\] which gives me \[\int\limits\limits_{\pi/4}^{3\pi/4}\sqrt{1+\cot^2(x)}dx\] and after using the Pythagorean identity: \[\int\limits\limits\limits_{\pi/4}^{3\pi/4}\csc(x)dx\] NOW THEN integrating it: \[-\ln|cscx+cotx| |_{\pi/4}^{3\pi/4}\] And this turns out to be 0. An arc-length of 0 is what throws me off, this can't be right...right?

Logic007:

\(L=-\ln|cscx+cotx| |_{\pi/4}^{3\pi/4}\) \(L=-\ln|csc(\frac{3}{4}\pi)+cot(\frac{1}{4}\pi)|-(-\ln|csc(\frac{3}{4}\pi)+cot(\frac{1}{4}\pi)|)\) \(L=-\ln|\sqrt{2}+(-1)|-(-\ln|\sqrt{2}+(+1)|)=-In|\sqrt{2}-1|+In|\sqrt{2}+1|\) From here,switch the two natural log and use the quotient property to combine them into a single log: \(L=In|\sqrt{2}+1|-In|\sqrt{2}-1|=In|\frac{(\sqrt{2}+1)}{(\sqrt{2}-1)}|\) Then rationalize the denominator by multiplying the conjugate and simplifying and use the power property of logs,and you will get: \(L=In|\frac{(\sqrt{2}+1)^2}{1}|=In|(\sqrt{2}+1)^2)|=2In|\sqrt{2}+1|\)

Logic007:

\(\approx1.76\)

mhchen:

Hmmm well..I'm confused about the transition from line 2 to line 3, where cot(pi/4) becomes +1 or -1.

Logic007:

Ops sorry 4 the confusion >.< , there is some typo error in line 2...

Logic007:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Logic007 \(L=-\ln|cscx+cotx| |_{\pi/4}^{3\pi/4}\) \(L=-\ln|csc(\frac{3}{4}\pi)+\color{blue}{cot(\frac{1}{4}\pi)}|-(-\ln|csc(\frac{3}{4}\pi)+\color{blue}{cot(\frac{1}{4}\pi)}|)\) \(L=-\ln|\sqrt{2}+(-1)|-(-\ln|\sqrt{2}+(+1)|)=-In|\sqrt{2}-1|+In|\sqrt{2}+1|\) From here,switch the two natural log and use the quotient property to combine them into a single log: \(L=In|\sqrt{2}+1|-In|\sqrt{2}-1|=In|\frac{(\sqrt{2}+1)}{(\sqrt{2}-1)}|\) Then rationalize the denominator by multiplying the conjugate and simplifying and use the power property of logs,and you will get: \(L=In|\frac{(\sqrt{2}+1)^2}{1}|=In|(\sqrt{2}+1)^2)|=2In|\sqrt{2}+1|\) \(\color{#0cbb34}{\text{End of Quote}}\) \(L=-\ln|csc(\frac{3}{4}\pi)+\color{red}{cot(\frac{3}{4}\pi)}|-(-\ln|csc(\frac{1}{4}\pi)+\color{red}{cot(\frac{1}{4}\pi)}|)\)

Logic007:

@mhchen

Logic007:

\(cot(\frac{\pi}{4})=+1\)

Logic007:

\(cot(\frac{3\pi}{4})=-1\)

mhchen:

omg you're right dam so that's what I missed on my midterm xD Thanks for clarifying cause half of my class told me they got 0 on that question and we all missed it lol

Logic007:

Haha,no prob. Thanks for asking. ^